### v1

@@ 4.2 Results
Results under outlier-aware TTA setting. We first evaluate TTA methods on the datasets with outliers. The results for CIFAR datasets are illustrated in Table 1. We observe that a simple baseline method, Tent [51], has a competitive generalization performance on distribution shift (achieving an accuracy of $80.1\%$ on CIFAR10-C). However, it exhibits substantial degradation in the AUC score compared to Source, for instance, dropping from $70.4\%$ to $48.7\%$ on CIFAR10-C with Noise outlier dataset. This decline could be attributed to Tent’s indiscriminate minimization of entropy across all samples, leading to a gradual loss of its ability to discern outliers. In addition, our experimental findings demonstrate that with judicious parameter selection, TTA methods can enhance not only the generalization capability against distribution shift but also the detection ability for outliers. Compared to existing TTA methods, STAMP achieves the most favorable performance on both classification and outlier detection. On CIFAR10-C, STAMP achieves the highest ACC, and $+4.1\%$ huge improvements in AUC compared to the previous methods. Regarding CIFAR100-C, STAMP exhibits $+2.8\%$ and $+5.7\%$ improvements compared to the most effective TTA methods on ACC and AUC, respectively. We also report the results on the ImageNet-C in Table 2. STAMP still achieves the highest performance. Specifically, we outperform existing TTA methods with $+1.2\%,+1.6\%$ on H-score with Places365-C and Textures-C outlier datasets, respectively.
@@ 4.3 Ablation Study
Furthermore, we conduct an ablation study on each component of our method, as summarized in Table 5. We observe that the effect of combining test-time step size decay and SAM is better than using either of them alone. Step size decay can improve the ACC but result in AUC decline. However, combining SAM and step size decay mitigates the decline in AUC. In addition, after the introduction of self-weighted entropy minimization, the AUC of the model further increased to $77.7\%$. Notably, by removing reliable class-balanced memory from our method, the ACC decreases from $63.7\%$ to $59.3\%$. It reveals that reliable class-balanced memory significantly contributes to enhancing the generalization ability to distribution shift. Moreover, by removing the self-weighted entropy minimization in STAMP, its AUC drops from $80.0\%$ to $75.0\%$. Self-weighted entropy minimization plays an important role in outlier detection. Besides decreasing the sample entropy, self-weighted entropy minimization incorporates entropy into the optimization weight, allowing outliers that deviate from the majority of samples to receive lower weights, consequently increasing their entropy. This mechanism enhances outlier detection performance.
@@ 0.B.1 Discussion about Augmentation-averaged Robust Estimation
0.B.2 Additional Ablation Study
Table 7: SAM ablation on CIFAR100-C with SVHN-C.
ACC AUC
STAMP-w/o-SAM 63.6 79.6
STAMP 63.7 80.0
Table 8: Memory ablation on ImageNet-C.
Places365-ACC Places365-AUC Textures-ACC Textures-AUC
STAMP-w/o-RBM 43.9 76.5 43.8 71.1
STAMP 45.6 76.8 46.8 71.4
We provide the additional ablation study of SAM optimizer and reliable class-balanced memory in Table 7 and Table 8.

### v2

@@ 4.2 Results
Results under outlier-aware TTA setting. We first evaluate TTA methods on the datasets with outliers. The results for CIFAR datasets are illustrated in Table 1. We observe that a simple baseline method, Tent [51], has a competitive generalization performance on distribution shift (achieving an accuracy of $80.1\%$ on CIFAR10-C). However, it exhibits substantial degradation in the AUC score compared to Source, for instance, dropping from $70.4\%$ to $48.7\%$ on CIFAR10-C with Noise outlier dataset. This decline could be attributed to Tent’s indiscriminate minimization of entropy across all samples, leading to a gradual loss of its ability to discern outliers. In addition, our experimental findings demonstrate that with judicious parameter selection, TTA methods can enhance not only the generalization capability against distribution shift but also the detection ability for outliers. Compared to existing TTA methods, STAMP achieves the most favorable performance on both classification and outlier detection. On CIFAR10-C, STAMP achieves the highest ACC, and $+5.1\%$ huge improvements in AUC compared to the previous methods. Regarding CIFAR100-C, STAMP exhibits $+2.8\%$ and $+7.3\%$ improvements compared to the most effective TTA methods on ACC and AUC, respectively. We also report the results on the ImageNet-C in Table 2. STAMP still achieves the highest performance. Specifically, we outperform existing TTA methods with $+1.6\%,+2.2\%$ on H-score with Places365-C and Textures-C outlier datasets, respectively.
@@ 4.3 Ablation Study
Furthermore, we conduct an ablation study on each component of our method, as summarized in Table 5. We observe that the effect of combining test-time step size decay and SAM is better than using either of them alone. Step size decay can improve the ACC but result in AUC decline. However, combining SAM and step size decay mitigates the decline in AUC. In addition, after the introduction of self-weighted entropy minimization, the AUC of the model further increased to $77.7\%$. Notably, by removing reliable class-balanced memory from our method, the ACC decreases from $63.7\%$ to $59.3\%$. It reveals that reliable class-balanced memory significantly contributes to enhancing the generalization ability to distribution shift. Moreover, by removing the self-weighted entropy minimization in STAMP, its AUC drops from $82.1\%$ to $75.1\%$. Self-weighted entropy minimization plays an important role in outlier detection. Besides decreasing the sample entropy, self-weighted entropy minimization incorporates entropy into the optimization weight, allowing outliers that deviate from the majority of samples to receive lower weights, consequently increasing their entropy. This mechanism enhances outlier detection performance.